OpenStudy (little.cherry23):

Help! i will fan and medal 1.)simplify the radical expression please show all of your steps sqrt363-3sqrt27

OpenStudy (michaelbp):

√363 - 3√27 = √(121*3) - 3√(9*3) = 11√3 - 3*3√3 = 2√3

OpenStudy (little.cherry23):

is it correct ?

OpenStudy (michaelbp):

im pretty sure

OpenStudy (michaelbp):

your welcome?

OpenStudy (little.cherry23):

sorry i was writting it down xD